OpenStudy (anonymous):

can someone help me find the quadratic equation? 1. maximum value 9; zeros of f are -6 and 0. i only know how to do it when given the x intercepts, not zero. thanks!

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=a\cdot (x-h)^2+k }\) that is a form of an opening down parabola, (where a is a real number, such that, a<0) with vertex (or maximum in our case) of (h,k).

OpenStudy (solomonzelman):

So far you know that k=9 (that is your maximum). \(\large\color{black}{ \displaystyle y=(-a)\cdot (x-h)^2+9 }\) and you know that (0,0) and (0,-6) are x-intercept

OpenStudy (solomonzelman):

So, since parabola is symmetric you can right of the bet conclude that it is centered at: Center x-value =(-6+0)/2 = -3 (at x=-3) So, \(\large\color{black}{ \displaystyle y=(-a)\cdot (x--3)^2+9 }\) \(\large\color{black}{ \displaystyle y=(-a)\cdot (x+3)^2+9 }\)

OpenStudy (solomonzelman):

oh, I wrote -a, so a should be greater than zero, for the coefficient -a to be neg.

OpenStudy (solomonzelman):

Use any of the points to find a. \(\large\color{black}{ \displaystyle 0=(-a)\cdot (0+3)^2+9 }\) and solve for a.

OpenStudy (anonymous):

ohhh ok. so whenever i am given zeros, i plug it into the first equation?

OpenStudy (anonymous):

a=1?

OpenStudy (solomonzelman):

Well, if you are given the zeros, you know the vertex x value. Because the parabola is always symmetric, so if it has zeros of x=c and x=d, then the x-value of the vertex is =(c+d)/2 or the midpoint X between the x-intercepts.

OpenStudy (solomonzelman):

-9=-9a a=1 precisely right!

OpenStudy (anonymous):

also, do i always have to set a to -a?

OpenStudy (solomonzelman):

Well, you can either say that it is -a, and solve for "a", but when you are writing you would write a negative a coefficient. \(\large\color{black}{ \displaystyle y=(-a)\cdot (x+3)^2+9 }\) you found a=1, so \(\large\color{black}{ \displaystyle y=(-\color{red}{1})\cdot (x+3)^2+9 }\)

OpenStudy (solomonzelman):

OR////

OpenStudy (solomonzelman):

Well, you can either say that it is a \(\large\color{black}{ \displaystyle y=(a)\cdot (x+3)^2+9 }\) then you would find that a=-1, and you would put it right in there \(\large\color{black}{ \displaystyle y=(\color{red}{-1})\cdot (x+3)^2+9 }\)

OpenStudy (solomonzelman):

You still have that -1 coefficient...

OpenStudy (anonymous):

Oh i see. THANK YOU SOO MUCH! You've been very helpful!!!

OpenStudy (solomonzelman):

Anytime... also do you know how I knew that my a is negative?

OpenStudy (solomonzelman):

I am saying, do you know how I figured that the coefficient in front of x^2 must be negative?

OpenStudy (anonymous):

Does it depend on whether the equation is minimum or maximum?

OpenStudy (solomonzelman):

Yes, basically very good. (Well if we ae taking about a vertically (up/down) opening parabola) So if it is going down, the coefficient is negative and if it is going up, the coefficient is positive

OpenStudy (solomonzelman):

So an opening down parab will have a maximum (but not min) and vice versa. opening up parab will have a minimum (but not max)

OpenStudy (anonymous):

I understand now. Thanks again!

OpenStudy (solomonzelman):

You are welcome!